OpenStudy (anonymous):

Square root property to solve equation: This one is a bit bigger: (3x+7)^2=8, show me how to do this please? :) The second part confuses me

OpenStudy (anonymous):

you need to use square root? you can't do \((a+b)^2=a^2+2ab+b^2\) ? and \(\sqrt8=\sqrt{2\cdot4}=2\sqrt2\)

OpenStudy (anonymous):

are you sure? wouldn't it just be sqrt2?

OpenStudy (anonymous):

but also you should remember that \(\sqrt{x^2}=|x|\) yes i am sure i showed you that 8=2*4 and you can square root out of 4 so it leaves \(2\sqrt2\)

OpenStudy (anonymous):

Oh! ^^ okay! i see now, thank you :)

OpenStudy (anonymous):

what answer you got?

OpenStudy (mathteacher1729):

The answer is not x = sqrt(2). See below for steps. (sorry, I'm too lazy to type them.) :-p